find the numerical value of y'(10) for y = sinx given that:

sin 0 = 0.000 , sin 10 = 0.1736 , sin 20 = 0.3420 , sin 30 = 0.5000 , sin 40 =0.6428


1
Expert's answer
2021-01-17T17:42:54-0500

The difference table is


"\begin{matrix}
 x & y & \Delta y & \Delta^2y & \Delta^3y & \Delta^4y \\
 0 & 0.0000 & & & & \\
 & &0.1736 & & & \\
10 & 0.1736 & & -0.0052 & & \\
 & &0.1684 & & -0.0052 & \\
20 & 0.3420 & & -0.0104 & & 0.0004 \\
 & & 0.1580 & & -0.0048 & \\
30 & 0.5000 & & -0.0152 & & \\
 & & 0.1428 & & & \\
40 & 0.6428 & & & & \\
\end{matrix}"

"h=\dfrac{10}{180}\pi=\dfrac{\pi}{18}"

"y'(10)=\dfrac{1}{h}(\Delta y_0-\dfrac{1}{2}\Delta^2y_0+\dfrac{1}{3}\Delta^3y_0-\dfrac{1}{4}\Delta^4y_0)"

"=\dfrac{18}{\pi}\bigg(0.1736-\dfrac{1}{2}(-0.0052)+\dfrac{1}{3}(-0.0052)"

"-\dfrac{1}{4}(0.0004)\bigg)\approx0.9990"

"y'(10)=0.9990"